OpenStudy (anonymous):

Keith want to create a drink that is 40% juice. How much of a 10% juice solution should he add to 100 milliliters of 100% grape juice to obtain the 40% mixture?

OpenStudy (anonymous):

@Chris182aa x = amount of 10% 100 = amount of 100% x + 100 = amount of 40% mix 10x + 100(100) = 40(x + 100) 10x + 10,000 = 40x + 4,000 -10x - 4000 - 10x - 4000 6000 = 30x answer -> x = 6000 ÷ by 30 = 200 mL

OpenStudy (anonymous):

i ment to put the divide symbol sorry

OpenStudy (nurali):

Let x = ml of 10% solution to be added 0.1x = ml of juice in 10% solution 10ml = ml of juice in 100% solution 0.1x+10/x+10=0.4 0.1x+10=0.4(x+10) 0.1x+10=0.4x+4 0.3x=6 x=20 20 ml of 10% solution needs to be added